One pipe fills a pool in 6 hours . A second pipe can fill the same pool in 4 hours, how long would it take to fill the pool if both pipes were working?
Answer by stanbon(48558)

One pipe fills a pool in 6 hours . A second pipe can fill the same pool in 4 hours, how long would it take to fill the pool if both pipes were working?
----------
1st pipe DATA:
Time = 6 hrs/job ; Rate = 1/6 job/hr.
------------------------------
2nd pipe DATA:
Time = 4 hrs/job ; Rate = 1/4 job/hr.
--------------------------------
Together DATA
Time = x hrs/job ; Rate = 1/x job/hr
------------------------
EQUATION:
rate + rate = together rate
1/6 + 1/4 = 1/x
2x + 3x = 12
5x = 12
x = 2.4 hrs (time to fill the pool using both pipes)
===============
Cheers,
Stan H.